# Fixed-sparsity matrix approximation from matrix-vector products

### v1

@@ 2 An algorithm and upper bound
$\operatorname{\mathbb{E}}\bigl{[}\|\mathbf{G}^{\dagger}\|_{\mathsf{F}}^{2}\bigr{]}=\frac{p}{p-q-1}.$ (5)
@@
The expression Equation 4 is an elementary calculation; see for instance [HMT11, Proposition A.1]. The expression Equation 5 follows from the fact that $\mathbf{G}^{\mathsf{T}}\mathbf{G}$ is invertible with probability one, and $(\mathbf{G}^{\mathsf{T}}\mathbf{G})^{-1}$ has a inverse Wishart distribution, which, for $p-q\geq 2$ has mean $\mathbf{I}/(p-q-1)$ [Mui82, §3.2 (12)]. Since $\|\mathbf{G}^{\dagger}\|_{\mathsf{F}}^{2}=\operatorname{tr}((\mathbf{G}^{\dagger})^{\mathsf{T}}(\mathbf{G}^{\dagger}))=\operatorname{tr}((\mathbf{G}^{\dagger})^{\mathsf{T}}(\mathbf{G}^{\dagger}))$, the result follows from the linearity of the expectation; see for instance [HMT11, Proposition A.5]. ∎
@@
Our bound in Corollary 1 has an unfavorable $O(1/\delta)$ dependence on the failure probability $\delta$. One could apply Markov’s inequality to each row and Hoeffding’s inequality to the sum to obtain a dependence $O(\log(n/\delta))$. However this has a dependence on the dimension $n$ which we would like to avoid. In Appendix B we show that one can apply a high-dimensional analog of the “median trick” to obtain an algorithm with a $O(\log(1/\delta))$ failure probability (without any dependence on the dimensions $n$ and $d$).
@@
Algorithm 1 requires solving $n$ least squares problems with a coefficient matrix of size $m\times s$. So, in addition to the application dependent cost of computing $\mathbf{Z}=\mathbf{A}\mathbf{G}$, its runtime is just $O(nms^{2})$. There are a number of practical improvements which can be made upon implementation. First, for many sparsity patterns, the matrices $\mathbf{G}_{i}$ and $\mathbf{G}_{i+1}$ differ only by a permutation and low-rank update. Thus, by downdating/updating appropriate quantities, the cost of solving all $n$ least-squares problems may be lower than $n$ times the cost of solving a single system. In addition, a posteriori variance estimates could also be obtained through Jack-knife type techniques [ET23].
@@ 3 A lower-bound for adaptive algorithms
This implies that for certain hard instances of Problems 1 and 2, adaptivity can only improve constants; it will not lead to improved rates. If $\mathbf{A}$ is known to have a particular structure, it is possible that adaptive algorithms may perform better than non-adaptive algorithms for these problems.
We note that the condition $\varepsilon<c$ is benign. In particular, if $C<c/2$, then $m\leq Cs/\varepsilon$ implies $m\leq s/2$, in which case one cannot expect to solve Problem 1, even in the zero error case, due to a parameter counting argument.
@@ 3.1 Key technical tools
Our hard distribution will be $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$, where $\mathbf{G}\sim\operatorname{Gaussian}(d)$. This is a special case of a so-called Wishart matrix. Our lower bound will make use of the fact that the conditional distribution of a Wishart matrix after a sequence of adaptive matrix-vector queries still looks like a slightly smaller transformed Wishart matrix [BHSW20, Lemma 3.4]. Similar hard input distributions have been used in a number of lower-bounds for matvec query tasks [SER18, BHSW20, JPWZ21, Che+23]. We believe other simple distributions such as $\mathbf{A}=\mathbf{G}$ or $\mathbf{A}=\mathbf{G}+\mathbf{G}^{\mathsf{T}}$ would also suffice to prove something like Theorem 2. We have chosen to use $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$ because it is symmetric, and it allows us to use a conceptually intuitive anti-concentration result based on the Berry–Esseen theorem.
The following is Lemma 3.4 from [BHSW20], restated to suit our needs:
@@
Suppose $\mathbf{G}\sim\operatorname{Gaussian}(d)$. Let $\mathbf{x}_{1},\ldots,\mathbf{x}_{t}$ and $\mathbf{y}_{1}=\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{x}_{1},\ldots,\mathbf{y}_{t}=\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{x}_{t}$ be such that, for each $j=1,\ldots,t$, $\mathbf{x}_{j}$ was chosen based only on the query vectors $\mathbf{x}_{1},\ldots,\mathbf{x}_{j-1}$ and the outputs $\mathbf{y}_{1},\ldots,\mathbf{y}_{j-1}$.
Then, there is is an $n\times n$ orthonormal matrix $\mathbf{V}_{t}$ and a $n\times n$ matrix $\mathbf{\Delta}_{t}$, each constructed solely as a function of $\mathbf{x}_{1},\ldots,\mathbf{x}_{t}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t}$, and a matrix $\mathbf{G}_{t}\sim\operatorname{Gaussian}(d-t)$ independent of $\mathbf{x}_{1},\ldots,\mathbf{x}_{t}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t}$ such that
$\mathbf{V}_{t}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{V}_{t}=\mathbf{\Delta}_{t}+\begin{bmatrix}\mathbf{0}_{t,t}&\mathbf{0}_{t,d-t}\\ \mathbf{0}_{d-t,t}&\mathbf{G}_{t}^{\mathsf{T}}\mathbf{G}_{t}\end{bmatrix}.$
We have included a proof in Section A.1 for completeness.
We will also use the following bound about the anti-concentration of independent random variables, which we prove in Section A.2. This is an immediate consequence of the Berry–Esseen Theorem and a basic anti-concentration result for Gaussians.
@@
Using Proposition 4, we can derive a more specific consequence which we will use directly in the proof of Theorem 2. The proof of this result is also contained in Section A.2.
@@ 3.2 Proof of Theorem 2
We now have the tools necessary to prove Theorem 2. The general strategy will be to show that the conditional distribution of a Wishart matrix after a sequence of (adaptive) queries is hard to approximate. That is, that the on-sparsity entries are anti-concentrated (conditioned on the queries) relative to the off-sparsity mass, which is $O(d^{3/2})$ with high probability.
Proof of Theorem 2.
Fix $\gamma\in(0,1)$. We will call values depending only on $\gamma$ “constants”, and do not track their dependence on $\gamma$ explicitly.
Let $C^{\prime}$ denote the absolute constant in Lemma 1. We will make the following assignments, which are listed in the order they appear: [^3] We have labeled the constants so that if $c_{i}$ depends on $c_{j}$, then $i>j$.
$c_{5}$ $=\frac{c_{4}(1-c_{1})c_{2}}{6c_{3}}$ $c_{6}$ $=\max\left\{2,\frac{C^{\prime}}{(1-c_{1})c_{5}\alpha^{2}c_{1}^{6}}\right\}$ $c_{1}$ $=\frac{\gamma}{4+2\gamma}$ $c_{4}$ $=\alpha^{2}c_{1}^{2}$
$c_{2}$ $=\frac{\gamma}{4+2\gamma}$ $c_{3}$ $=150$ $C_{1}$ $=c_{5}c_{1}$ $\alpha$ $=1/100.$
$c$ $=\min\{1/c_{6},c_{5}/2\}$ $C$ $=2C_{1}.$
Fix $\varepsilon>0$. We will show that if
$m\leq C_{1}\frac{s}{\varepsilon},\qquad\frac{c_{5}}{\varepsilon}\in\mathbb{Z},\qquad\varepsilon<\frac{1}{c_{6}}\leq\frac{1}{2},$ (7)
The assumption $c_{5}/\varepsilon\in\mathbb{Z}$ will subsequently be removed.
With this in mind, set
$d=c_{5}\frac{s}{\varepsilon},\quad\text{suppose}\quad m\leq C_{1}\frac{s}{\varepsilon}=c_{1}d,\quad\text{and define}\quad k=d-m.$
Let $\widehat{\mathbf{G}}\sim\operatorname{Gaussian}(d)$ and define $\mathbf{A}=\widehat{\mathbf{G}}^{\mathsf{T}}\widehat{\mathbf{G}}$. Suppose we do $m$ adaptive queries to $\mathbf{A}$. Proposition 3 implies that there exists an $d\times d$ matrix $\mathbf{\Delta}$, and $d\times k$ matrix $\mathbf{V}$ with orthonormal columns, both constructed solely as a function of the queries and measurements, and a matrix $\mathbf{G}\sim\operatorname{Gaussian}(k)$ independent of the queries and measurements such that
$\mathbf{A}=\mathbf{\Delta}+\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}},\qquad\mathbf{W}=\mathbf{G}^{\mathsf{T}}\mathbf{G}.$
@@
$P=\Big{\{}(i,j)\in[d]\times[d]:\big{|}[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}-[\mathbf{T}]_{i,j}\big{|}^{2}>c_{4}k\Big{\}},$
and the events
$E=\Big{\{}|P|\geq c_{2}ds\Big{\}},\qquad F=\Big{\{}\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\leq c_{3}d^{3}\Big{\}}.$
@@
$\|\mathbf{S}\circ\mathbf{A}-\mathbf{T}\|_{\mathsf{F}}^{2}\geq\sum_{(i,j)\in P}\big{|}[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}-[\mathbf{T}]_{i,j}\big{|}^{2}\geq|P|\cdot c_{4}k\geq c_{2}ds\cdot c_{4}k.$
Since $m\leq c_{1}d$, we have that $k\geq(1-c_{1})d$. By definition of $d$, $s=d\varepsilon/c_{5}$. By definition, $c_{5}=c_{4}(1-c_{1})c_{2}/(6c_{3})$. We then have that
$\|\mathbf{S}\circ\mathbf{A}-\mathbf{T}\|_{\mathsf{F}}^{2}\geq c_{2}d\cdot\frac{\varepsilon d}{c_{5}}\cdot c_{4}\cdot(1-c_{1})d=6\varepsilon c_{3}d^{3}.$
Finally, if $F$ holds, then
$\|\mathbf{S}\circ\mathbf{A}-\mathbf{T}\|_{\mathsf{F}}^{2}\geq 6\varepsilon\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}.$
Now noting that $\mathbf{S}\circ\mathbf{A}$ and $\mathbf{T}$ each have disjoint support from $\mathbf{A}-\mathbf{S}\circ\mathbf{A}$, and that $(1+6\varepsilon)\geq(1+2\varepsilon)^{2}$ for all $\varepsilon\in(0,1/2)$,
$\|\mathbf{A}-\mathbf{T}\|_{\mathsf{F}}^{2}\geq(1+6\varepsilon)\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\geq(1+2\varepsilon)^{2}\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}.$
Thus, under the assumptions of Equation 7, it suffices to show that $E$ and $F$ hold simultaneously with probability at least $24/25$. Recall our assumption on $m$:
$m\leq c_{1}d=c_{5}c_{1}\cdot\frac{s}{\varepsilon}.$
It is easy to show that $\operatorname{\mathbb{E}}[\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}]\leq\operatorname{\mathbb{E}}[\|\mathbf{A}\|_{\mathsf{F}}^{2}]\leq 3d^{3}$ (see Fact 2 for a derivation), so since $c_{3}=150$, an application of Markov’s inequality implies $F$ happens with probability at least $49/50$. Thus, if we show that $E$ also happens with probability at least $49/50$, then the result follows by a union bound.
It remains to analyze the probability of $E$. Towards this end, recall $c_{1}=\gamma/(4+2\gamma)$, let $\mathbf{v}_{i}$ be the $i$th row of $\mathbf{V}$, and let
$L=\big{\{}i\in[d]:\|\mathbf{v}_{i}\|_{2}^{2}\geq c_{1}\big{\}}.$
@@
We must have $|L|\geq 2d/(2+\gamma)$. Otherwise, since $\mathbf{V}$ has orthonormal columns so that $\|\mathbf{v}_{i}\|_{2}^{2}\leq 1$, we would have:
$k=\|\mathbf{V}\|_{\mathsf{F}}^{2}=\sum_{i=1}^{d}\|\mathbf{v}_{i}\|_{2}^{2}<\sum_{i\not\in L}c_{1}+\sum_{i\in L}1<c_{1}d+\frac{2d}{2+\gamma}=(1-c_{1})d,$
which contradicts our assumption $m\leq c_{1}d\Leftrightarrow k\geq(1-c_{1})d$.
Now, define
$M=\big{\{}(i,j):i,j\in L~{}\text{and}~{}[\mathbf{S}]_{i,j}=1\big{\}}.$
Recall $c_{2}=\gamma/(4+2\gamma)=c_{1}$. Since $|L|\geq 2d/(2+\gamma)$, Lemma 2 gives that $|M|\geq\gamma ds/(2+\gamma)=2c_{2}ds$.
We claim it suffices to show that
$\forall(i,j)\in M:\operatorname{\mathbb{P}}\bigl{[}(i,j)\in P\bigr{]}\geq\frac{99}{100}.$ (9)
Indeed, by Markov’s inequality, we would immediately have that
$\operatorname{\mathbb{P}}\Biggl{[}\sum_{(i,j)\in M}\mathds{1}[(i,j)\not\in P]\leq\frac{100}{2}\cdot\frac{|M|}{100}\Biggr{]}\geq 1-\frac{1}{50}=\frac{49}{50}.$
Since $|M|\geq 2c_{2}ds$, this then implies that
$\operatorname{\mathbb{P}}\bigl{[}E\bigr{]}=\operatorname{\mathbb{P}}\biggl{[}|P|\geq c_{2}ds\biggr{]}\geq\operatorname{\mathbb{P}}\biggl{[}|P|\geq\frac{|M|}{2}\biggr{]}\geq\frac{49}{50}.$
@@
We must show Equation 9. Fix arbitrary $(i,j)\in M$. Since $[\mathbf{S}]_{i,j}=1$, note that $[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}=\mathbf{x}_{i}^{\mathsf{T}}\mathbf{x}_{j}$, where $\mathbf{x}_{i}=\mathbf{G}\mathbf{v}_{i}$ and $\mathbf{x}_{j}=\mathbf{G}\mathbf{v}_{j}$ and $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ are the $i$-th and $j$-th rows of $\mathbf{V}$ respectively. Since $i,j\in L$, we have that $\|\mathbf{v}_{i}\|_{2}^{2},\|\mathbf{v}_{j}\|_{2}^{2}\geq c_{1}$. With this in mind, recall that by the assumptions of Equation 7, $1/\epsilon\geq c_{6}$, where $c_{6}\geq C^{\prime}/((1-c_{1})c_{5}\alpha^{2}c_{1}^{6})$ by definition. Then
$k=d-m\geq(1-c_{1})\cdot c_{5}\frac{s}{\varepsilon}\geq(1-c_{1})c_{5}c_{6}\geq\frac{C^{\prime}}{\alpha^{2}c_{1}^{6}}\geq\frac{C^{\prime}}{\alpha^{2}\|\mathbf{v}_{i}\|_{2}^{6}\|\mathbf{v}_{j}\|_{2}^{6}}.$
Hence, applying Lemma 1 and the definition $c_{4}=\alpha^{2}c_{1}^{2}$,
@@
$\operatorname{\mathbb{P}}\Bigl{[}(i,j)\not\in P\Bigr{]}$ $=\operatorname{\mathbb{P}}\Bigl{[}\big{|}[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}-[\mathbf{T}]_{i,j}\big{|}<\sqrt{c_{4}k}\Bigr{]}$

### v3

@@ 2 An algorithm and upper bound
$\operatorname{\mathbb{E}}\bigl{[}\|\mathbf{G}^{\dagger}\|_{\mathsf{F}}^{2}\bigr{]}=\frac{q}{p-q-1}.$ (5)
@@
The expression Equation 4 is an elementary calculation; see for instance [halko_martinsson_tropp_11, Proposition A.1]. The expression Equation 5 follows from the fact that $\mathbf{G}^{\mathsf{T}}\mathbf{G}$ is invertible with probability one, and $(\mathbf{G}^{\mathsf{T}}\mathbf{G})^{-1}$ has a inverse Wishart distribution, which, for $p-q\geq 2$ has mean $\mathbf{I}/(p-q-1)$ [muirhead_82, §3.2 (12)]. Since $\|\mathbf{G}^{\dagger}\|_{\mathsf{F}}^{2}=\operatorname{tr}((\mathbf{G}^{\dagger})^{\mathsf{T}}(\mathbf{G}^{\dagger}))=\operatorname{tr}((\mathbf{G}^{\dagger})^{\mathsf{T}}(\mathbf{G}^{\dagger}))$, the result follows from the linearity of the expectation; see for instance [halko_martinsson_tropp_11, Proposition A.5]. ∎
@@
Our bound in Corollary 1 has an unfavorable $O(1/\delta)$ dependence on the failure probability $\delta$. One could apply Markov’s inequality to each row and Hoeffding’s inequality to the sum to obtain a dependence $O(\log(n/\delta))$. However this has a dependence on the dimension $n$ which we would like to avoid. In Appendix C we show that one can apply a high-dimensional analog of the “median trick” to obtain an algorithm with a $O(\log(1/\delta))$ failure probability (without any dependence on the dimensions $n$ and $d$).
@@
Algorithm 1 requires solving $n$ least squares problems with a coefficient matrix of size $m\times s$. So, in addition to the application dependent cost of computing $\mathbf{Z}=\mathbf{A}\mathbf{G}$, its runtime is just $O(nms^{2})$. There are a number of practical improvements which can be made upon implementation. First, for many sparsity patterns, the matrices $\mathbf{G}_{i}$ and $\mathbf{G}_{i+1}$ differ only by a permutation and low-rank update. Thus, by downdating/updating appropriate quantities, the cost of solving all $n$ least-squares problems may be lower than $n$ times the cost of solving a single system. In addition, a posteriori variance estimates could also be obtained through Jack-knife type techniques [epperly_tropp_23].
@@ 3 A lower-bound for adaptive algorithms
This implies that for certain hard instances of Problems 1 and 2, our Corollary 1 and thus Theorem 1 are optimal up to constants. In particular, adaptivity can only improve constants; it will not lead to an improved dependence on $s$ or $\epsilon$. If $\mathbf{A}$ is known to have a particular structure, it is possible that adaptive algorithms may perform better than non-adaptive algorithms for these problems.
We note that the condition $\varepsilon<c$ is benign. In particular, if $C<c/2$, then $m\leq Cs/\varepsilon$ implies $m\leq s/2$, in which case one cannot solve Problem 1, even in the zero error case, due to a parameter counting argument.
@@ 3.1 Key technical tools
Our hard distribution will be $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$, where $\mathbf{G}\sim\operatorname{Gaussian}(d)$. This is a special case of a so-called Wishart matrix. Our lower bound will make use of the fact that the conditional distribution of a Wishart matrix after a sequence of adaptive matrix-vector queries still looks like a slightly smaller transformed Wishart matrix [braverman_hazan_simchowitz_woodworth_20, Lemma 3.4]. Similar hard input distributions have been used in a number of lower-bounds for matvec query tasks [simchowitz_elalaoui_recht_18, braverman_hazan_simchowitz_woodworth_20, jiang_pham_woodruff_zhang_21, chewi_dediospont_li_lu_narayanan_23]. We believe other simple distributions such as $\mathbf{A}=\mathbf{G}$ or $\mathbf{A}=\mathbf{G}+\mathbf{G}^{\mathsf{T}}$ would also suffice to prove something like Theorem 2. We have chosen to use $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$ because it is symmetric, and it allows us to use a conceptually intuitive anti-concentration result based on the Berry–Esseen theorem.
The following is essentially Lemma 3.4 from [braverman_hazan_simchowitz_woodworth_20], restated to suit our needs:
@@
Suppose $\mathbf{G}\sim\operatorname{Gaussian}(d,r)$. Let $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}$ and $\mathbf{y}_{1}=\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{x}_{1},\ldots,\mathbf{y}_{m}=\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{x}_{m}$ be such that, for each $j=1,\ldots,m$, $\mathbf{x}_{j}$ was chosen based only on the query vectors $\mathbf{x}_{1},\ldots,\mathbf{x}_{j-1}$ and the outputs $\mathbf{y}_{1},\ldots,\mathbf{y}_{j-1}$.
Then, there is an $n\times n$ orthonormal matrix $\mathbf{V}_{m}$ and an $n\times n$ matrix $\mathbf{\Delta}_{m}$, each constructed solely as functions of $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}$, and a matrix $\mathbf{G}_{m}\sim\operatorname{Gaussian}(d-m,r-m)$ independent of $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}$ such that
$\mathbf{V}_{m}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{V}_{m}=\mathbf{\Delta}_{m}+\begin{bmatrix}\mathbf{0}_{m,m}&\mathbf{0}_{m,d-m}\\ \mathbf{0}_{d-m,m}&\mathbf{G}_{m}^{\mathsf{T}}\mathbf{G}_{m}\end{bmatrix}.$
We have included a proof in Section B.1 for completeness.
We will also use the following bound about the anti-concentration of independent random variables, which we prove in Section B.2. This is an immediate consequence of the Berry–Esseen Theorem and a basic anti-concentration result for Gaussians.
@@
Using Proposition 4, we can derive a more specific consequence which we will use directly in the proof of Theorem 2. The proof of this result is also contained in Section B.2.
@@ 3.2 Proof of Theorem 2
We now have the tools necessary to prove Theorem 2. The general strategy will be to show that the conditional distribution of a Wishart matrix after a sequence of (adaptive) queries is hard to approximate. That is, that the on-sparsity entries are anti-concentrated (conditioned on the queries) relative to the off-sparsity mass, which is $O(d^{3/2})$ with high probability. We will then use this result to prove Theorem 2.
Lemma 3 (Main technical lemma).
Fix $\gamma\in(0,1)$. Then there exist constants $c_{1},c_{2},c_{3}>0$ (depending only on $\gamma$) such that the following holds:
Suppose $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$ where $\mathbf{G}\sim\operatorname{Gaussian}(r,d)$. Then, for any sparsity pattern $\mathbf{S}$ whose rows and columns each have between $\gamma s$ and $s$ nonzero entries, and for any (possibly randomized) algorithm that uses $m$ (possibly adaptive) matrix-vector queries to $\mathbf{A}$ to output $\widetilde{\mathbf{A}}$ with $\widetilde{\mathbf{A}}\circ\mathbf{S}=\widetilde{\mathbf{A}}$, if $m\leq c_{1}d$ and, for any $\alpha\in(0,1)$, $r>m+c_{3}/\alpha^{2}$, then
$\operatorname{\mathbb{P}}\Big{[}\|\mathbf{S}\circ\mathbf{A}-\widetilde{\mathbf{A}}\|_{\mathsf{F}}^{2}<c_{2}\alpha^{2}ds(r-m)\Big{]}<\alpha.$
Proof.
Fix $\gamma\in(0,1)$. Let $C$ denote the absolute constant in Lemma 1, define
$c_{1}=\frac{\gamma}{4+2\gamma}\qquad c_{2}=\frac{c_{1}^{2}}{4}\qquad c_{3}=\frac{C}{c_{1}^{6}},$
and suppose $m$, $r$, and $d$ are integers such that
$m\leq c_{1}d,\qquad r-m>\frac{c_{3}}{\alpha^{2}}.$
Suppose we do $m$ (possibly adaptive) queries to $\mathbf{A}$. Proposition 3 implies that there exists an $d\times d$ matrix $\mathbf{\Delta}$, and $d\times(d-m)$ matrix $\mathbf{V}$ with orthonormal columns, both constructed solely as functions of the queries and measurements, and a matrix $\widehat{\mathbf{G}}\sim\operatorname{Gaussian}(r-m,d-m)$ independent of the queries and measurements such that
$\mathbf{A}=\mathbf{\Delta}+\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}},\qquad\mathbf{W}=\widehat{\mathbf{G}}^{\mathsf{T}}\widehat{\mathbf{G}}.$
@@
$P=\Big{\{}(i,j)\in[d]\times[d]:\big{|}[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}-[\mathbf{T}]_{i,j}\big{|}^{2}>(\alpha/2)^{2}c_{1}(r-m)\Big{\}},$
and the event
$E=\Big{\{}|P|\geq c_{1}ds\Big{\}}.$
@@
$\|\mathbf{S}\circ\mathbf{A}-\mathbf{T}\|_{\mathsf{F}}^{2}$ $\geq\sum_{(i,j)\in P}\big{|}[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}-[\mathbf{T}]_{i,j}\big{|}^{2}$
$\geq|P|\cdot(\alpha/2)^{2}c_{1}(r-m)\geq c_{1}ds\cdot(\alpha/2)^{2}c_{1}(r-m)=\alpha^{2}c_{2}ds(r-m),$
so it remains to show $\operatorname{\mathbb{P}}[E]\geq 1-\alpha$.
Towards this end, let $\mathbf{v}_{i}$ be the $i$th row of $\mathbf{V}$, and define
$I=\big{\{}i\in[d]:\|\mathbf{v}_{i}\|_{2}^{2}\geq c_{1}\big{\}},\qquad M=\big{\{}(i,j)\in I\times I:[\mathbf{S}]_{i,j}=1\big{\}}.$
@@
We must have $|I|\geq 2d/(2+\gamma)$. Otherwise, since $\mathbf{V}$ has orthonormal columns so that $\|\mathbf{v}_{i}\|_{2}^{2}\leq 1$, we would have:
$d-m=\|\mathbf{V}\|_{\mathsf{F}}^{2}=\sum_{i=1}^{d}\|\mathbf{v}_{i}\|_{2}^{2}<\sum_{i\in I^{\mathsf{c}}}c_{1}+\sum_{i\in I}1<c_{1}d+\frac{2d}{2+\gamma}=(1-c_{1})d,$
which contradicts our assumption $m\leq c_{1}d\Leftrightarrow d-m\geq(1-c_{1})d$. Then, since $|I|\geq 2d/(2+\gamma)$, Lemma 2 and our assumption on $\mathbf{S}$ give that
$|M|\geq\frac{\gamma ds}{2+\gamma}=2c_{1}ds.$
@@
We now show that if $(i,j)\in M$ then $(i,j)\in P$ with high probability. Fix arbitrary $(i,j)\in M$ (and note that this means $i,j\in I$). Since $[\mathbf{S}]_{i,j}=1$, note that $[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}=\mathbf{x}_{i}^{\mathsf{T}}\mathbf{x}_{j}$, where $\mathbf{x}_{i}=\widehat{\mathbf{G}}\mathbf{v}_{i}$ and $\mathbf{x}_{j}=\widehat{\mathbf{G}}\mathbf{v}_{j}$ and $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ are the $i$-th and $j$-th rows of $\mathbf{V}$ respectively. We also have that $\|\mathbf{v}_{i}\|_{2}^{2},\|\mathbf{v}_{j}\|_{2}^{2}\geq c_{1}$. With this in mind, our choice of constants implies
$r-m\geq\frac{c_{3}}{\alpha^{2}}=\frac{C}{\alpha^{2}c_{1}^{6}}\geq\frac{C}{\alpha^{2}\|\mathbf{v}_{i}\|_{2}^{6}\|\mathbf{v}_{j}\|_{2}^{6}}.$
Hence, applying Lemma 1,
@@
$\operatorname{\mathbb{P}}\Bigl{[}(i,j)\not\in P\Bigr{]}$ $=\operatorname{\mathbb{P}}\Bigl{[}\big{|}[\mathbf{S}\circ\mathbf{V}\mathbf{W}\mathbf{V}^{\mathsf{T}}]_{i,j}-[\mathbf{T}]_{i,j}\big{|}<(\alpha/2)c_{1}\sqrt{r-m}\Bigr{]}$